OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (owlcoffee):

\[-5(x+1)=10\] let's apply distributive: \[-5x-5=10\] if we add 5 to both sides we get: \[-5x=15\] You can take over from here on :)

OpenStudy (anonymous):

@Owlcoffee the brackets are absolute brackets

OpenStudy (austinl):

That would only be one answer. You would need to solve for the other answer. However, because this is multiple choice it narrows down the correct answer.

OpenStudy (owlcoffee):

oh my. then we'll have to evaluate that absolute value, making the opposite of that~

OpenStudy (anonymous):

So B??

OpenStudy (anonymous):

how do you make the opposite of it ?

OpenStudy (anonymous):

−5|x + 1| = 10 first step?

OpenStudy (anonymous):

divide -5 ?

OpenStudy (anonymous):

| x + 1| = -2

OpenStudy (owlcoffee):

When we solve for an abs. value, we get two ecuations: 1) \[-5(x+1)=10\] 2)\[-5(-\left[ x+1 \right])=10\] Solve for x on those two and you'll get the answer.

OpenStudy (anonymous):

thanks

OpenStudy (austinl):

\(−5|x + 1| = 10\) Divide by -5. \(|x+1|=-2\) That means that we have, \(x+1=-2\) & \(-(x+1)=-2\) \(-x-1=-2\) Solve each of them, \(x=-3\) \(x=1\) Now this is what I get, however wolfram alpha says no solutions.

OpenStudy (anonymous):

the issue is that the absolute brackets makes everything positive when the other side of the equation has a negative sign I guess thats why theres no solution - the sign on the LHS and RHS always conflict